OpenStudy (anonymous):

If the radius of a circle is multiplied by 4, then how many times larger is the circumference of the new circle?

OpenStudy (unklerhaukus):

\[A_{\text{circle}}=\pi r^2\]

OpenStudy (anonymous):

would it be 4

OpenStudy (anonymous):

these are the given choices 2. 4 8 and pie

OpenStudy (unklerhaukus):

\[A_{\text{circle}_1}=\pi (r_1)^2\] \[A_{\text{circle}_2}=\pi (r_2)^2\] \[r_2=4r_1\]

OpenStudy (anonymous):

so 4 would be correct

OpenStudy (unklerhaukus):

ops looks like i have mis read the question sorry @alijahe

OpenStudy (unklerhaukus):

the question is asking us to compair the Circumference of the circles , not the areas.

OpenStudy (unklerhaukus):

\[C=2 \pi r\]\[C_1=2 \pi r_1\]\[C_2=2 \pi r_2\] \[r_2=4r_1\]

OpenStudy (anonymous):

ohh ok

OpenStudy (unklerhaukus):

\[\frac{C_2}{ C_1}=\frac{2\pi r_2}{2 \pi r_1}=\frac{2\pi \times4 r_1}{2 \pi r_1}\]

OpenStudy (unklerhaukus):

what do you get ?

OpenStudy (anonymous):

4 after i complete all the work

OpenStudy (unklerhaukus):

that is correct.

OpenStudy (anonymous):

thank you